Assume an integer is represented on 1 byte. What is the problem with adding the following two binary numbers?

0111 1111+

0100 0001

--------------

1100 0000

A. There is nothing wrong

B. The answer is correct; it should be 0100 00100

C. The answer is incorrect; it should be 0110 0000

D. The sum of the two numbers is to big to be represented only one byte.
Answer #1

Answer: Correct option is A.here is nothing wrong

Explanation:

rules for binary addition:

i) 0 + 0 = 0

ii) 1 + 0 = 1

iii) 1 + 1 = 10 (binary for 2)

iv) 1 + 1 + 1 = 11 (binary for 3)

we have,

01111111

+  01000001

---------------------

11000000

Addition is correct as per rule of binary addition so nothing is wrong.